# Private Heterogeneous Federated Learning Without a Trusted Server Revisited: Error-Optimal and Communication-Efficient Algorithms for Convex Losses

### v2

@@ Abstract
We revisit the problem of federated learning (FL) with private data from people who do not trust the server or other silos/clients. In this context, every silo (e.g. hospital) has data from several people (e.g. patients) and needs to protect the privacy of each person’s data (e.g. health records), even if the server and/or other silos try to uncover this data. Inter-Silo Record-Level Differential Privacy (ISRL-DP) prevents each silo’s data from being leaked, by requiring that silo $i$’s communications satisfy item-level differential privacy. Prior work (Lowy and Razaviyayn, 2023b) characterized the optimal excess risk bounds for ISRL-DP algorithms with homogeneous (i.i.d.) silo data and convex loss functions. However, two important questions were left open: (1) Can the same excess risk bounds be achieved with heterogeneous (non-i.i.d.) silo data? (2) Can the optimal risk bounds be achieved with fewer communication rounds? In this paper, we give positive answers to both questions. We provide novel ISRL-DP FL algorithms that achieve the optimal excess risk bounds in the presence of heterogeneous silo data. Moreover, our algorithms are more communication-efficient than the prior state-of-the-art. For smooth loss functions, our algorithm achieves the optimal excess risk bound and has communication complexity that matches the non-private lower bound. Additionally, our algorithms are more computationally efficient than the previous state-of-the-art.
@@ 1 Introduction
Following Lowy et al. (2023a); Lowy and Razaviyayn (2023b); Liu et al. (2022) (among others), this work considers inter-silo record-level DP (ISRL-DP). ISRL-DP requires that the full transcript of messages sent by silo $i$ satisfy item-level DP, for all silos $i$. Thus, ISRL-DP guarantees the privacy of each silo’s local data, even in the presence of an adversary with access to the server, other silos, or the communication channels between silos and the server. See Figure 1 for an illustration. If each silo only has one record, then ISRL-DP reduces to local DP (Kasiviswanathan et al., 2011; Duchi et al., 2013). However, in the FL setting where each silo has many records, ISRL-DP FL is more appropriate and permits much higher accuracy than local DP (Lowy and Razaviyayn, 2023b). Moreover, ISRL-DP implies user-level DP if the ISRL-DP parameters are sufficiently small (Lowy and Razaviyayn, 2023b). Thus, ISRL-DP is a practical privacy notion for cross-silo and cross-device FL when individuals do not trust the server or other silos/clients/devices with their sensitive data.
@@
The quality of a private algorithm $\mathcal{A}$ for solving Problem FL is measured by its excess risk (a.k.a. excess population/test loss), $\mathbb{E}[F(\mathcal{A}(\mathbf{X}))]-F^{*}$, where $F^{*}=\inf_{w\in\mathcal{W}}F(w)$ and the expectation is over the random draw of the data $\mathbf{X}=(X_{1},\ldots,X_{N})$ as well as the randomness of $\mathcal{A}$Lowy and Razaviyayn (2023b) characterized the minimax optimal excess risk (up to logarithms) for convex ISRL-DP FL in the homogeneous (i.i.d.) case:
@@
where $d$ is the dimension of the parameter space and $\varepsilon,\delta$ are the privacy parameters. For heterogeneous FL, the state-of-the-art excess risk bound is $O\left(1/\sqrt{Nn}+\left(\sqrt{d\log(1/\delta)}/(\varepsilon n\sqrt{N})\right)^{4/5}\right)$, assuming smoothness of the loss function (Lowy and Razaviyayn, 2023b). Lowy and Razaviyayn (2023b) left the following question open:
@@
In practical FL settings (especially cross-device settings), it can be expensive or slow for silos to communicate with the server or with each other (Kairouz et al., 2021). Thus, another important desiderata for FL algorithms is communication efficiency, which is measured by communication complexity: the number of communication rounds $R$ that are required to achieve excess risk $\alpha$. In the homogeneous setting, the state-of-the-art communication complexity for an algorithm achieving optimal excess risk is due to Lowy and Razaviyayn (2023b, Theorem D.1):
@@
For non-smooth loss functions, our communication complexity is
$R_{\text{non-smooth}}=\widetilde{O}\left(\min\left\{(Nn)^{1/2},\frac{N^{1/2}n\varepsilon}{d^{1/2}}\right\}\right),$ (5)
@@
In the homogeneous case, the state-of-the-art gradient complexity for an ISRL-DP FL algorithm that attains optimal excess risk is due to Lowy and Razaviyayn (2023b, Theorem D.1). For smooth losses and $\varepsilon=\Theta(1)$, Lowy and Razaviyayn (2023b) obtained the follow gradient complexity
@@
In Theorem 3.4, we also improve over the state-of-the-art subgradient complexity bound of Lowy and Razaviyayn (2023b, Theorem D.2) for non-smooth losses. Moreover, in contrast to these earlier results, we do not assume homogeneous data.
@@
Table 1: Comparison vs. SOTA for Smooth Loss Functions. [LR’23] refers to Lowy and Razaviyayn (2023b); we omit logs and fix $M=N$, $d=\Theta(n)$, $\varepsilon=\Theta(1)$, $L=\beta=D=1$. See theorems and the appendix for more general cases.
@@
Table 2: Comparison vs. SOTA for Nonsmooth Loss Functions. First 3 algorithms use Nesterov smoothing.
@@
[LR’23] Alg. 2 (i.i.d.) optimal $Nn$ $N^{2}n^{2}$
[LR’23] Alg. 1 (non-i.i.d.) suboptimal $N^{1/3}n^{1/3}$ $Nn$
Alg. 1 (non-i.i.d.) optimal $N^{1/2}n^{1/2}$ Omitted (see Remark 3.3)
@@
Our ISRL-DP Algorithm 1 for smooth FL builds on the ISRL-DP Accelerated MB-SGD approach used by Lowy and Razaviyayn (2023b) for federated empirical risk minimization (ERM). We call this algorithm repeatedly to iteratively solve a carefully chosen sequence of regularized ERM subproblems, using the localization technique of Feldman et al. (2020). We obtain our novel optimal excess risk bounds for heterogeneous FL with an algorithmic stability (Bousquet and Elisseeff, 2002) argument. The key observation is that the stability and generalization of regularized ERM in Shalev-Shwartz et al. (2009) does not require the data to be identically distributed. By combining this observation with a bound for ISRL-DP Accelerated MB-SGD, we obtain our excess risk bound.
It is not immediately clear that the approach just described should work, because the iterates of Accelerated MB-SGD are not necessarily stable. The instability of acceleration may explain why Lowy and Razaviyayn (2023b) used ISRL-DP Accelerated MB-SGD only for ERM and not for minimizing the population risk. We overcome this obstacle with an alternative analysis that leverages the stability of regularized ERM, the convergence of ISRL-DP Accelerated MB-SGD to an approximate minimizer of the regularized empirical loss, and localization. This argument enables us to show that our algorithm has optimal excess risk. By carefully choosing algorithmic parameters, we also obtain state-of-the-art communication and gradient complexities.
To extend our algorithm to the non-smooth case, we use two different techniques. One is Nesterov smoothing (Nesterov, 2005), which results in an algorithm with favorable communication complexity. The other approach is to replace the ISRL-DP Accelerated MB-SGD ERM subsolver by an ISRL-DP Subgradient Method. This technique yields an algorithm with favorable (sub)gradient complexity. Third, we use randomized convolution smoothing, as in Kulkarni et al. (2021), to obtain another favorable gradient complexity bound.
@@ 2 Localized ISRL-DP Accelerated MB-SGD for Smooth Losses
We start with the smooth case. Combining iterative localization techniques of Feldman et al. (2020); Asi et al. (2021) with the multi-stage ISRL-DP Accelerated MB-SGD of Lowy and Razaviyayn (2023b), our proposed Algorithm 1 achieves optimal excess risk and state-of-the-art communication complexity for heterogeneous FL under ISRL-DP.
@@
Algorithm 2 Accelerated ISRL-DP MB-SGD (Lowy and Razaviyayn, 2023b)
@@
We recall ISRL-DP Accelerated MB-SGD in Algorithm 2. It is a distributed, ISRL-DP version of the AC-SA algorithm (Ghadimi and Lan, 2012). For strongly convex losses, the multi-stage implementation of ISRL-DP Accelerated MB-SGD, given in Algorithm 5 in Appendix B, offers improved excess risk (Lowy and Razaviyayn, 2023b). Algorithm 5 is a distributed, ISRL-DP version of the multi-stage AC-SA (Ghadimi and Lan, 2013).
@@
The excess risk bound in (9) matches the optimal i.i.d risk bound up to log factors (cf. Theorem 2.2 in Lowy and Razaviyayn (2023b)), even when silo data is arbitrarily heterogeneous across silos. To the best of our knowledge, our algorithm is the first to have this property, resolving an open question of Lowy and Razaviyayn (2023b). The prior state-of-the-art bound in Lowy and Razaviyayn (2023b, Theorem 3.1) is suboptimal by a factor of $\widetilde{O}((\sqrt{d}/(\varepsilon n\sqrt{N}))^{1/5})$.
@@
Theorem 2.1 trivially extends to the unconstrained optimization setting in which $D=\|w_{0}-w^{*}\|$ for $w^{*}\in\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}F(w)$. Moreover, our excess risk bound is still optimal for the unconstrained case: a matching lower bound is obtained by combining the technique for unconstrained DP lower bounds in Liu and Lu (2021) with the constrained ISRL-DP FL lower bounds of Lowy and Razaviyayn (2023b).
@@
$R$ $\gtrsim N^{1/4}\left(\min\left\{\sqrt{n},\frac{\varepsilon n}{\sqrt{d\ln(1/\delta)}}\right\}\right)^{1/2}$
×$\;\;\;\;\times\min\left(\frac{\sqrt{\beta D}}{\sqrt{L}},\frac{\zeta_{*}}{\sqrt{\beta LD}}\right).$
@@
where $\hat{w}_{i}=\operatorname*{arg\,min}_{w\in\mathcal{W}}\hat{F}_{i}(w)$ for $i\in[\tau]$. Then, it remains to bound (10) and (11), respectively. To this end, we first show that $w_{i}$ in Algorithm 1 is close to $\hat{w}_{i}$ for all $i\in[\tau]$. Here, we mainly leverage the excess empirical risk bound of Algorithm 5 for ERM with strongly convex and smooth loss functions. In particular, by the empirical risk bound in Lowy and Razaviyayn (2023b, Theorem F.1) and $\lambda_{i}$-strongly convex of $\hat{F}_{i}$, we can show that the following bound holds for all $i\in[\tau]$:
@@
3 Error-Optimal Heterogeneous ISRL-DP FL for Non-Smooth Losses
In this section, we turn to the case of non-smooth loss functions. We modify Algorithm 1 to obtain two algorithms that can handle the non-smooth case. Our algorithms are the first to achieve the optimal excess risk for heterogeneous ISRL-DP FL with non-smooth loss functions.
3.1 Error-Optimal and Communication-Efficient ISRL-DP FL for Non-Smooth Losses
Our first algorithm is based on the technique of Nesterov smoothing (Nesterov, 2005): For a non-smooth function $f$, Moreau-Yosida regularization is used to approximate $f$ by the $\beta$-smooth $\beta$-Moreau envelope
@@ 3.1 Error-Optimal and Communication-Efficient ISRL-DP FL for Non-Smooth Losses
Theorem 3.1 (Non-smooth FL via smoothing).
@@
Theorem 3.1 gives the first optimal rate for non-smooth heterogeneous FL under ISRL-DP. Note that the current best result for heterogeneous FL in Lowy and Razaviyayn (2023b) is suboptimal and holds only for the smooth case. One can combine the same smoothing technique above with the algorithm in Lowy and Razaviyayn (2023b) to obtain a risk bound of $O(1/\sqrt{nN}+(\sqrt{d\ln(1/\delta)}/{\varepsilon n\sqrt{N}})^{2/3})$ for the non-smooth case, which is again suboptimal.
Remark 3.3 (Improved communication complexity).
The communication complexity bound in Theorem 3.1 improves over the previous state-of-the-art result for an algorithm achieving optimal excess risk (Lowy and Razaviyayn, 2023b); recall (3). Moreover, the result of Lowy and Razaviyayn (2023b) assumed i.i.d. silo data, whereas our result holds for the non-i.i.d. case.
We have omitted the gradient complexity for this smoothing approach. This is mainly because the gradient computation of $f_{\beta}$ needs additional computation in the form of the prox operator. One may consider using an approximate prox operator to get a handle on the total gradient complexity as in Bassily et al. (2019). However, this approach would still introduce an additional $\Theta(n^{3})$ gradient complexity per step. Thus, a natural question is whether we can design a more computation-efficient algorithm for the non-smooth case — providing a motivation for our next algorithm.
3.2 Error-Optimal and Computationally Efficient ISRL-DP FL for Non-Smooth Losses
We propose another variation of Algorithm 1 that uses subgradients to handle the non-smooth case in a computationally efficient way: see Algorithm 4. Algorithm 4 follows the same structure as Algorithm 1: we iteratively solve a carefully chosen sequence of regularized ERM problems with a ISRL-DP solver and use localization. Compared to Algorithm 1 for the smooth case, Algorithm 4 does not use an accelerated solver (due to non-smoothness). Instead, we use ISRL-DP Minibatch Subgradient method (Algorithm 3) to solve the non-smooth strongly convex ERM problem in each phase of Algorithm 4. There are two key differences between our subroutine Algorithm 3 and the ISRL-DP MB-SGD of Lowy and Razaviyayn (2023b): (i) Instead of the gradient, a subgradient of the non-smooth objective is used in 6; (ii) A different and simpler averaging step in 11 is used for strongly convex non-smooth losses.

### v3

@@ Abstract
We revisit the problem of federated learning (FL) with private data from people who do not trust the server or other silos/clients. In this context, every silo (e.g. hospital) has data from several people (e.g. patients) and needs to protect the privacy of each person’s data (e.g. health records), even if the server and/or other silos try to uncover this data. Inter-Silo Record-Level Differential Privacy (ISRL-DP) prevents each silo’s data from being leaked, by requiring that silo $i$’s communications satisfy item-level differential privacy. Prior work (Lowy and Razaviyayn, 2023a) characterized the optimal excess risk bounds for ISRL-DP algorithms with homogeneous (i.i.d.) silo data and convex loss functions. However, two important questions were left open: (1) Can the same excess risk bounds be achieved with heterogeneous (non-i.i.d.) silo data? (2) Can the optimal risk bounds be achieved with fewer communication rounds? In this paper, we give positive answers to both questions. We provide novel ISRL-DP FL algorithms that achieve the optimal excess risk bounds in the presence of heterogeneous silo data. Moreover, our algorithms are more communication-efficient than the prior state-of-the-art. For smooth loss functions, our algorithm achieves the optimal excess risk bound and has communication complexity that matches the non-private lower bound. Additionally, our algorithms are more computationally efficient than the previous state-of-the-art.
@@ 1 Introduction
Following Lowy and Razaviyayn (2023a); Lowy et al. (2023a); Liu et al. (2022) (among others), this work considers inter-silo record-level DP (ISRL-DP). ISRL-DP requires that the full transcript of messages sent by silo $i$ satisfy item-level DP, for all silos $i$. Thus, ISRL-DP guarantees the privacy of each silo’s local data, even in the presence of an adversary with access to the server, other silos, or the communication channels between silos and the server. See Figure 1 for an illustration. If each silo only has one record, then ISRL-DP reduces to local DP (Kasiviswanathan et al., 2011; Duchi et al., 2013). However, in the FL setting where each silo has many records, ISRL-DP FL is more appropriate and permits much higher accuracy than local DP (Lowy et al., 2023a). Moreover, ISRL-DP implies user-level DP if the ISRL-DP parameters are sufficiently small (Lowy et al., 2023a). Thus, ISRL-DP is a practical privacy notion for cross-silo and cross-device FL when individuals do not trust the server or other silos/clients/devices with their sensitive data.
@@
The quality of a private algorithm $\mathcal{A}$ for solving Problem FL is measured by its excess risk (a.k.a. excess population/test loss), $\mathbb{E}[F(\mathcal{A}(\mathbf{X}))]-F^{*}$, where $F^{*}=\inf_{w\in\mathcal{W}}F(w)$ and the expectation is over the random draw of the data $\mathbf{X}=(X_{1},\ldots,X_{N})$ as well as the randomness of $\mathcal{A}$Lowy et al. (2023a) characterized the minimax optimal excess risk (up to logarithms) for convex ISRL-DP FL in the homogeneous (i.i.d.) case:
@@
where $d$ is the dimension of the parameter space and $\varepsilon,\delta$ are the privacy parameters. For heterogeneous FL, the state-of-the-art excess risk bound is $O\left(1/\sqrt{Nn}+\left(\sqrt{d\log(1/\delta)}/(\varepsilon n\sqrt{N})\right)^{4/5}\right)$, assuming smoothness of the loss function (Lowy et al., 2023a). Lowy et al. (2023a) left the following question open:
@@
In practical FL settings (especially cross-device settings), it can be expensive or slow for silos to communicate with the server or with each other (Kairouz et al., 2021). Thus, another important desiderata for FL algorithms is communication efficiency, which is measured by communication complexity: the number of communication rounds $R$ that are required to achieve excess risk $\alpha$. In the homogeneous setting, the state-of-the-art communication complexity for an algorithm achieving optimal excess risk is due to Lowy et al. (2023a, Theorem D.1):
@@
For nonsmooth loss functions, our communication complexity is
$R_{\text{nonsmooth}}=\widetilde{O}\left(\min\left\{(Nn)^{1/2},\frac{N^{1/2}n\varepsilon}{d^{1/2}}\right\}\right),$ (5)
@@
In the homogeneous case, the state-of-the-art gradient complexity for an ISRL-DP FL algorithm that attains optimal excess risk is due to Lowy et al. (2023a, Theorem D.1). For smooth losses and $\varepsilon=\Theta(1)$, Lowy et al. (2023a) obtained the follow gradient complexity
@@
In Theorem 3.5, we also improve over the state-of-the-art subgradient complexity bound of Lowy et al. (2023a, Theorem D.2) for nonsmooth losses. Moreover, in contrast to these earlier results, we do not assume homogeneous data.
@@
Table 1: Comparison vs. SOTA for Smooth Loss Functions. [LR’23] refers to Lowy et al. (2023a); we omit logs and fix $M=N$, $d=\Theta(n)$, $\varepsilon=\Theta(1)$, $L=\beta=D=1$. See theorems and the appendix for more general cases.
@@
Table 2: Comparison vs. SOTA for Nonsmooth Loss Functions. First 3 algorithms use Nesterov smoothing and we omit the gradient complexity results for them (see Theorem 3.1).
@@
[LR’23] Alg. 2 (i.i.d.) optimal $Nn$ Omitted
[LR’23] Alg. 1 (non-i.i.d.) suboptimal $N^{1/3}n^{1/3}$ Omitted
Alg. 1 (non-i.i.d.) optimal $N^{1/2}n^{1/2}$ Omitted
@@
Our ISRL-DP Algorithm 1 for smooth FL builds on the ISRL-DP Accelerated MB-SGD approach used by Lowy et al. (2023a) for federated empirical risk minimization (ERM). We call this algorithm repeatedly to iteratively solve a carefully chosen sequence of regularized ERM subproblems, using the localization technique of Feldman et al. (2020). We obtain our novel optimal excess risk bounds for heterogeneous FL with an algorithmic stability (Bousquet and Elisseeff, 2002) argument. The key observation is that the stability and generalization of regularized ERM in Shalev-Shwartz et al. (2009) does not require the data to be identically distributed. By combining this observation with a bound for ISRL-DP Accelerated MB-SGD, we obtain our excess risk bound.
It is not immediately clear that the approach just described should work, because the iterates of Accelerated MB-SGD are not necessarily stable. The instability of acceleration may explain why Lowy et al. (2023a) used ISRL-DP Accelerated MB-SGD only for ERM and not for minimizing the population risk. We overcome this obstacle with an alternative analysis that leverages the stability of regularized ERM, the convergence of ISRL-DP Accelerated MB-SGD to an approximate minimizer of the regularized empirical loss, and localization. This argument enables us to show that our algorithm has optimal excess risk. By carefully choosing algorithmic parameters, we also obtain state-of-the-art communication and gradient complexities.
To extend our algorithm to the nonsmooth case, we use two different techniques. One is Nesterov smoothing (Nesterov, 2005), which results in an algorithm with favorable communication complexity. The other approach is to replace the ISRL-DP Accelerated MB-SGD ERM subsolver by an ISRL-DP Subgradient Method. This technique yields an algorithm with favorable (sub)gradient complexity. Third, we use randomized convolution smoothing, as in Kulkarni et al. (2021), to obtain another favorable gradient complexity bound.
@@ 2 Localized ISRL-DP Accelerated MB-SGD for Smooth Losses
We start with the smooth case. Combining iterative localization techniques of Feldman et al. (2020); Asi et al. (2021) with the multi-stage ISRL-DP Accelerated MB-SGD of Lowy et al. (2023a), our proposed Algorithm 1 achieves optimal excess risk and state-of-the-art communication complexity for heterogeneous FL under ISRL-DP.
@@
Algorithm 2 Accelerated ISRL-DP MB-SGD (Lowy et al., 2023a)
@@
We recall ISRL-DP Accelerated MB-SGD in Algorithm 2. It is a distributed, ISRL-DP version of the AC-SA algorithm (Ghadimi and Lan, 2012). For strongly convex losses, the multi-stage implementation of ISRL-DP Accelerated MB-SGD, given in Algorithm 5 in Appendix B, offers improved excess risk (Lowy et al., 2023a). Algorithm 5 is a distributed, ISRL-DP version of the multi-stage AC-SA (Ghadimi and Lan, 2013).
@@
The excess risk bound in (9) matches the optimal i.i.d risk bound up to log factors (cf. Theorem 2.2 in Lowy et al. (2023a)), even when silo data is arbitrarily heterogeneous across silos. To the best of our knowledge, our algorithm is the first to have this property, resolving an open question of Lowy et al. (2023a). The prior state-of-the-art bound in Lowy et al. (2023a, Theorem 3.1) is suboptimal by a factor of $\widetilde{O}((\sqrt{d}/(\varepsilon n\sqrt{N}))^{1/5})$.
@@
Theorem 2.1 trivially extends to the unconstrained optimization setting in which $D=\|w_{0}-w^{*}\|$ for $w^{*}\in\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}F(w)$. Moreover, our excess risk bound is still optimal for the unconstrained case: a matching lower bound is obtained by combining the technique for unconstrained DP lower bounds in Liu and Lu (2021) with the constrained ISRL-DP FL lower bounds of Lowy et al. (2023a).
@@
$R$ $\gtrsim N^{1/4}\left(\min\left\{\sqrt{n},\frac{\varepsilon n}{\sqrt{d\ln(1/\delta)}}\right\}\right)^{1/2}\min\left(\frac{\sqrt{\beta D}}{\sqrt{L}},\frac{\zeta_{*}}{\sqrt{\beta LD}}\right).$
@@
where $\hat{w}_{i}=\operatorname*{arg\,min}_{w\in\mathcal{W}}\hat{F}_{i}(w)$ for $i\in[\tau]$. Then, it remains to bound (10) and (11), respectively. To this end, we first show that $w_{i}$ in Algorithm 1 is close to $\hat{w}_{i}$ for all $i\in[\tau]$. Here, we mainly leverage the excess empirical risk bound of Algorithm 5 for ERM with strongly convex and smooth loss functions. In particular, by the empirical risk bound in Lowy et al. (2023a, Theorem F.1) and $\lambda_{i}$-strongly convex of $\hat{F}_{i}$, we can show that the following bound holds for all $i\in[\tau]$:
@@
3 Error-Optimal Heterogeneous ISRL-DP FL for Nonsmooth Losses
In this section, we turn to the case of nonsmooth loss functions. We modify Algorithm 1 to obtain algorithms that can handle nonsmooth losses. Our algorithms are the first to achieve the optimal excess risk for heterogeneous ISRL-DP FL with nonsmooth loss functions.
3.1 Reductions to the Smooth Case via Smoothing
Nesterov Smoothing:
Our first algorithm is based on the technique of Nesterov smoothing (Nesterov, 2005): For a nonsmooth function $f$, Moreau-Yosida regularization is used to approximate $f$ by the $\beta$-smooth $\beta$-Moreau envelope
@@ 3.1 Reductions to the Smooth Case via Smoothing
Theorem 3.1 (Nonsmooth FL via Nesterov smoothing).
@@
We have omitted the gradient complexity for this smoothing approach. This is mainly because the gradient computation of $f_{\beta}$ needs additional computation in the form of the prox operator. One may consider using an approximate prox operator to get a handle on the total gradient complexity as in Bassily et al. (2019). However, this approach would still introduce an additional $\Theta(n^{3})$ gradient complexity per step.
Convolutional Smoothing
An alternative approach is to use convolutional smoothing, in which we approximate $f(w,x_{i})$ by the smooth function $\mathbb{E}_{v\sim\mathcal{U}_{s}}f(w+v,x_{i})$, where $\mathcal{U}_{s}$ denotes the uniform distribution over the centered $\ell_{2}$ ball of radius $s$. We then apply Algorithm 1 to the smooth approximation of $f$ with slight changes. We defer details to Appendix D and only state simplfied results below.
Theorem 3.2 (Nonsmooth FL via convolutional smoothing).
Let $M=N$. Then, there exist parameter choices such that combining Algorithm 1 with convolutional smoothing (see Algorithm 6) yields an $(\varepsilon,\delta)$-ISRL-DP algorithm with optimal excess population risk as in (9). The communication complexity (in expectation) is
$\widetilde{O}\left(\max\left\{1,d^{1/4}\sqrt{M}\min\left\{\sqrt{n},\frac{\varepsilon n}{\sqrt{d\ln(1/\delta)}}\right\},\frac{\varepsilon^{2}n}{d\ln(1/\delta)}\right\}\right).$ (13)
When $d=\Theta(n)$, and $\varepsilon=\Theta(1)$, the gradient complexity is
$\widetilde{O}\left(M^{3/2}n^{3/4}+M^{5/4}n^{11/8}\right).$
Compared with Nesterov smoothing, convolutional smoothing approach has better gradient complexity with slightly worse communication complexity.
@@
By combining Algorithm 1 with Nesterov smoothing or convolutional smoothing, we give the first optimal rate for nonsmooth heterogeneous FL under ISRL-DP. Note that the previous best result for heterogeneous FL in Lowy et al. (2023a) is suboptimal and holds only for the smooth case. One can combine the same smoothing technique above with the algorithm in Lowy et al. (2023a) to obtain a risk bound of $O(1/\sqrt{nN}+(\sqrt{d\ln(1/\delta)}/{(\varepsilon n\sqrt{N})})^{2/3})$ for the nonsmooth case, which is again suboptimal.
Remark 3.4 (Improved communication complexity and gradient complexity).
The communication complexity of both smoothing approaches improves over the previous state-of-the-art result for an algorithm achieving optimal excess risk (Lowy et al., 2023a); recall (3). The convolutional smoothing approach also improves the gradient complexity. Moreover, the result of Lowy et al. (2023a) assumed i.i.d. silo data, whereas our result holds for the non-i.i.d. case.
Depending on the FL application, communication or computation may be more of a bottleneck. If communication efficiency is more of a priority, the Nesterov smoothing approach should be used. On the other hand, if computational efficiency is more pressing, convolutional smoothing is recommended.
3.2 A Direct Subgradient Algorithm
We propose another variation of Algorithm 1 that uses subgradients to handle the nonsmooth case in a direct and computationally efficient way: see Algorithm 4. Algorithm 4 follows the same structure as Algorithm 1: we iteratively solve a carefully chosen sequence of regularized ERM problems with a ISRL-DP solver and use localization. Compared to Algorithm 1 for the smooth case, Algorithm 4 does not use an accelerated solver (due to nonsmoothness). Instead, we use ISRL-DP Minibatch Subgradient method (Algorithm 3) to solve the nonsmooth strongly convex ERM problem in each phase of Algorithm 4. There are two key differences between our subroutine Algorithm 3 and the ISRL-DP MB-SGD of Lowy et al. (2023a): (i) Instead of the gradient, a subgradient of the nonsmooth objective is used in Line 5; (ii) A different and simpler averaging step in Line 10 is used for strongly convex nonsmooth losses. We state simplified results below; the complete version and the proof can be found in Theorem G.1.